Verify that the given family of functions solves the differential equation.
(i) dy⁄dx= (1-2t)y², y=⅟(C – t + t²)
(ii) dy⁄dx= y² sin t, y=⅟(C + cos t)
1
Expert's answer
2020-07-02T19:09:21-0400

"(i) 
\frac{dy}{dt}=(1-2t)y^2\\[1 em]
\text{We separate y and t to get}\\[1 em]
\frac{dy}{y^2}=(1-2t)dt\\[1 em]
\text{Integrate}\\[1 em]
 
\int \frac{dy}{y^2}=\int(1-2t)dt\\[1 em]
\therefore \frac{-1}{y}=(t-t^2+c)\\[1 em]
\therefore y= \frac{-1}{t-t^2+c}=\frac{1}{C-t+t^2}\\[1 em]

(ii) 
 \frac{dy}{dt}=y^2 \sin t\\[1 em]
\text{We separate y and t to get}\\[1 em]
\frac{dy}{y^2}=\sin t ~dt \\[1 em]
\text{Integrate}\\[1 em]

\int \frac{dy}{y^2}=\sin t ~dt \\[1 em]
\therefore \frac{-1}{y}=-\ cos~ t +c\\[1 em]
\therefore y= \frac{-1}{-\ cos ~t +c}=\frac{1}{\ cos ~t +~C}\\[1 em]"